# VITA-1.5: Towards GPT-4o Level Real-Time Vision and Speech Interaction

### v3

@@ Abstract
Recent Multimodal Large Language Models (MLLMs) have typically focused on integrating visual and textual modalities, with less emphasis placed on the role of speech in enhancing interaction. However, speech plays a crucial role in multimodal dialogue systems, and implementing high-performance in both vision and speech tasks remains a significant challenge due to the fundamental modality differences. In this paper, we propose a carefully designed multi-stage training methodology that progressively trains LLM to understand both visual and speech information, ultimately enabling fluent vision and speech interaction. Our approach not only preserves strong vision-language capacity, but also enables efficient speech-to-speech dialogue capabilities without separate ASR and TTS modules, significantly accelerating multimodal end-to-end response speed. By comparing our method against state-of-the-art counterparts across benchmarks for image, video, and speech tasks, we demonstrate that our model is equipped with both strong visual and speech capabilities, making near real-time vision and speech interaction. The training and inference codes have been released at https://github.com/VITA-MLLM/VITA (2K+ Stars by now).
@@ 1 Introduction
Recent advancements in MLLMs [13, 31, 67, 10, 49, 61, 42, 17] have led to significant progress, particularly in the integration of visual and textual modalities. The introduction of visual information into LLMs has notably enhanced model capabilities across a range of multimodal tasks. However, with the growing appeal of human-computer interaction, the role of the speech modality has become increasingly prominent, especially in the multimodal dialogue system. In such a system, speech not only serves as a key medium for information transmission but also greatly improves the naturalness and convenience of interactions. Consequently, integrating both visual and speech modalities to achieve high-performance multimodal interactions has emerged as a critical research focus.
The integration of vision and speech in MLLMs is not straightforward due to their inherently differences [40]. For example, visual data, such as images, convey spatial information, while speech data convey dynamic changes in time series. These fundamental differences pose challenges for simultaneous optimization of both modalities, often leading to conflicts during training. For instance, the inclusion of speech data may degrade performance on vision tasks, and vice versa. In addition, traditional speech-to-speech systems rely on separate modules for Automatic Speech Recognition (ASR) and Text-to-Speech, which can increase latency and reduce coherence, limiting their practicality in real-time applications [44, 16, 63].
@@ 2 Related Work
Recently, thanks to the rapid development of language models such as GPTs [41, 3], LLaMA [52, 53], Alpaca [48], Vicuna [12], and Mistral [24], researchers have successfully extended text comprehension to multimodal understanding/reasoning through techniques like multimodal alignment and instruction tuning. For example, models such as LLaVA [31], Qwen-VL [2], Cambrian-1 [51], Mini-Gemini [28], MiniCPM-V 2.5 [23], DeepSeek-VL [36], and SliME [66] have made significant advances in image perception and reasoning, while models like LongVA [65] and Video-LLaVA [29] have showcased the latest progress in video understanding. These models are increasingly capable of handling diverse data types, driving the continuous improvement of multimodal perception and understanding capabilities.
However, compared to proprietary models that support multiple modalities, including audio, image, and text (e.g., GPT-4o [42] and Gemini-Pro 1.5 [50]), most open-source models have primarily focused on image and text modalities [61]. Moreover, few open-source models have involved multimodal interaction capabilities, which is a relatively unexplored area. While works like VITA-1.0 [16] have made initial attempts to introduce speech for human-computer interaction, introducing additional speech data poses challenges to the model’s original multimodal abilities. Furthermore, speech generation typically relies on existing TTS systems, which often results in high latency, thus impacting user experience. In this paper, we present VITA-1.5 that leverages a refined training strategies, excelling in perceiving data across four modalities (video, image, text, and audio), while also realizing near real-time vision and speech interaction.
@@
Figure 2: Overall Architecture of VITA-1.5. The input side consists of vision and audio encoders, along with their adapters connected to a LLM. The output side has an end-to-end speech generation module, rather than directly using an external TTS model as the initial VITA-1.0 version [16].
@@ 3.1 Model Architecture
The overall architecture of VITA-1.5 is depicted in Fig. 2. The input side is the same as that of the VITA-1.0 version [16], that is, adopting the configuration of “Multimodal Encoder-Adaptor-LLM”. It combines the Vision/Audio Transformer and the Multi-Layer Connector with an LLM for joint training, aiming to enhance the unified understanding of vision, language, and audio. With respect to the output side, VITA-1.5 has its own end-to-end speech module, instead of using the external TTS model like the original VITA-1.0 version.
@@ 3.1.1 Visional Modality
Visual Encoder. VITA-1.5 adopts InternViT-300M [^1] https://huggingface.co/OpenGVLab/InternViT-300M-448px as the visual encoder, with an input image size of 448×448 pixels, generating 256 visual tokens per image. For high-resolution images, VITA-1.5 employs a dynamic patching [10] strategy to capture local details, improving the accuracy of image understanding.
@@ 3.1.2 Audio Modality
Speech Encoder. Similar to [56], our audio encoding module consists of multiple downsampling convolutional layers (4x downsampling) and 24 Transformer blocks (with a hidden size of 1024). The downsampling layers help reduce the frame rate of the audio features, improving the processing speed of LLM. The audio encoder has about 350M parameters and an output frame rate of 12.5Hz. Mel-filter bank features are used as the input of the audio encoder, with a window size of 25ms and a shift of 10ms [56].
@@
Speech Decoder. TiCodec [45] is used as our codec model, customizing a single codebook with a size of 1024. This single-codebook design simplifies the decoding process during the inference phase. The codec model is responsible for encoding continuous speech signals into discrete speech tokens with the frequency of 40Hz, and at the same time has the ability to decode them back into speech signals with the sample rate of 24, [^0] Hz.
The current LLM can only output text tokens, and the speech generation capability requires the LLM to be able to output speech tokens. To this end, we add two speech decoders after the text tokens following [56]: 1) Non-Autoregressive (NAR) Speech Decoder, which processes text tokens globally and models semantic features, with the aim of generating an initial distribution of speech tokens; 2) Autoregressive (AR) Speech Decoder generates higher quality speech tokens step by step, based on the speech information produced by the NAR decoder. The final sequence of speech tokens is then decoded into a continuous speech signal flow (waveform) using the speech decoder of the Codec model. We adopt 4 LLaMA decoder layers for both NAR and AR speech decoders, where the hidden size is 896 and the parameter size is about 120M.
@@ 3.2 Training Data
  • Image Captioning Data. Datasets such as ShareGPT4V [9], ALLaVA-Caption [6], SharedGPT4o-Image [^2] https://sharegpt4o.github.io/, and synthetic data are used to train the model to generate descriptive languages for images.
  • Image QA Data. Datasets like LLaVA-150K [^3] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K, LLaVA-Mixture-sample [31], LVIS-Instruct [55], ScienceQA [38], ChatQA [35], and subsets sampled from LLaVA-OV [26], such as general image QA and mathematical reasoning datasets, are utilized to train the model in answering image-based questions and performing visual reasoning tasks.
  • OCR & Diagram Data. This category supports the model in understanding OCR and diagram content, using datasets such as Anyword-3M [54], ICDAR2019-LSVT [^4] http://icdar2019.org/, UReader [58], SynDOG555naver-clova-ix/synthdog-en, ICDAR2019-LSVT-QA [^6] http://icdar2019.org/, and corresponding data sampled from LLaVA-OV.
  • Video Data. Datasets like ShareGemini [47] and synthetic data are used to train the model to handle video inputs and perform tasks such as captioning and video-based QA.
  • Pure Text Data. This category enhances the model’s capability to understand and generate languages, facilitating text-based QA tasks.
@@
Table 1: Training data of multimodal instruction tuning. The images of the synthetic data come from open-source datasets like Wukong [19], LAION [46], and CC12M [5].
@@ 3.3.2 Stage 2: Audio Input Tuning
Stage 2.1 Audio Alignment. After completing the training of Stage 1, the model has developed a strong foundation in image and video understanding. In this stage, our goal is to reduce the discrepancy between audio and language based on Stage 1, enabling the LLM to understand audio inputs. The training data consists of 11,000 hours of speech-transcription pairs. We follow a two-step approach: (a) Speech Encoder Training: We adopt a training framework used in common speech recognition systems, using a Connectionist Temporal Classification (CTC) loss function [18] to train the speech encoder. The aim is for the encoder to predict the transcription text from the speech input. This step ensures that the audio encoder can extract speech features and map them to the text representation space. (b) Speech Adapter Training: After training the speech encoder, we integrate it with the LLM, using an audio adapter to introduce audio features into the input layer of the LLM. The training objective at this stage is to enable the LLM to output the transcription text of the speech data.
@@ 3.3.3 Stage 3: Audio Output Tuning
In the first two stages of training, the VITA-1.5 model has effectively developed its multimodal understanding capabilities. However, a crucial capacity, i.e., speech output, remains absent, which is essential for its role as an interactive assistant. To introduce speech output functionality without compromising the model’s fundamental abilities, we draw on the strategy [56], using 3,000 hours of text-speech data and employing a two-step training approach (see Fig. 3).
@@ 4.1 Vision-Language Evaluation
Baselines. We compare a series of open-source MLLMs, including VILA-1.5 [30], LLaVA-Next [25], CogVLM2 [22], InternLM-XComposer2.5 [64], Cambrian-1 [51], MiniCPM-V-2.6 [23], Ovis1.5 [39], InternVL-Chat-1.5, InternVL-2 [11], LLaVA-OV [26], and Video-LLaVA [29], SliME [66], and LongVA [65], as well as 5 closed-source MLLMs, including GPT-4V [^7] https://openai.com/index/gpt-4v-system-card/, GPT-4o [^8] https://openai.com/index/hello-gpt-4o/, GPT-4o-mini, Gemini 1.5 Pro [50], and Claude 3.5 Sonnet [^9] https://www.anthropic.com/news/claude-3-5-sonnet.
Evaluation Benchmarks. To assess the image perception and understanding capabilities of VITA-1.5, we utilize several evaluation benchmarks, including MME [14], MMBench [32], MMStar [8], MMMU [60], MathVista [37], HallusionBench [20], AI2D [21], OCRBench [34], and MMVet [59]. These benchmarks cover a wide range of aspects, including general multimodal capabilities (e.g., MME, MMBench, and MMMU), mathematical reasoning (MathVista), hallucination detection (HallusionBench), chart (AI2D) and OCR (OCRBench) understanding, providing a comprehensive evaluation results. For video understanding, we use representative evaluation benchmarks including Video-MME [15], MVBench [27], and TempCompass [33].
@@
Mini-Omini2 - - - 4.8 9.8 4.7 9.4
Freeze-Omini 2.8 12.6 14.2 4.2 10.2 4.1 10.5
Ours
@@ 4.2 Speech Evaluation
Baselines. The following three baseline models are used for comparison: Wav2vec2-base [1], Mini-Omini2 [57], Freeze-Omini [56], and VITA-1.0 [16].
Evaluation Benchmarks. The Mandarin Evaluation Sets consists of three datasets: aishell-1 [4], test net [7], and test meeting [62]. These datasets are used to evaluate the model’s performance on Mandarin speech. The evaluation metric is the Character Error Rate (CER). The English Evaluation Sets include four datasets: dev-clean, dev-other, test-clean, and test-other [43], which are used to evaluate the model’s performance on English speech. The evaluation metric is Word Error Rate (WER).
ASR Performance. The evaluation results in Table. 4 indicate that VITA-1.5 achieves leading accuracy in both Mandarin and English ASR tasks. This demonstrates that VITA-1.5 has successfully integrated advanced speech capability to support multimodal interaction.
5 Conclusion
In this paper, we has presented VITA-1.5, a multimodal LLM designed to integrate vision and speech through a carefully crafted three stage training strategy. By relieving the inherent conflicts between modalities, VITA-1.5 achieves robust capabilities in both vision and speech understanding, enabling efficient speech-to-speech interactions without relying on separate ASR or TTS modules. Extensive evaluations demonstrate that VITA-1.5 performs competitively across multimodal benchmarks. We hope that VITA-1.5 can take over the banner of VITA-1.0 and continue to promote the progress of open-source models in the field of real-time multimodal interaction.

### v4

@@ Abstract
Recent Multimodal Large Language Models (MLLMs) have typically focused on integrating visual and textual modalities, with less emphasis placed on the role of speech in enhancing interaction. However, speech plays a crucial role in multimodal dialogue systems, and implementing high-performance in both vision and speech tasks remains a challenge due to the fundamental modality differences. In this paper, we propose a carefully designed multi-stage training methodology that progressively trains LLM to understand both visual and speech information, ultimately enabling fluent vision and speech interaction. Our approach not only preserves strong vision-language capacity, but also enables efficient speech-to-speech dialogue capabilities without separate ASR and TTS modules, significantly accelerating multimodal end-to-end response speed. By comparing against state-of-the-art counterparts across benchmarks for image, video, and speech, we demonstrate that our omni model is equipped with both strong visual and speech capabilities, making omni understanding and interaction.
@@ 1 Introduction
Recent advancements in MLLMs (liu2023visual, ; zhan2024anygpt, ; zhang2023video, ; zhang2025speechact, ; sun2024video, ; shu2023audio, ; wen2025dycrowd, ; tang2025empowering, ; tu2024overview, ) have led to significant progress, particularly in integration of visual and textual modalities. The introduction of visual information into LLMs has notably enhanced model capabilities across various multimodal tasks. However, with the growing appeal of human-computer interaction, the role of the speech modality has become increasingly prominent, especially in multimodal dialogue systems. In such a system, speech not only serves as a key medium for information transmission but also greatly improves the naturalness and convenience of interactions. Consequently, integrating visual and speech modalities to achieve multimodal interactions has emerged as a critical research focus.
The integration of vision and speech in MLLMs is not straightforward due to their inherently differences (oneațua2022improving, ). For example, visual data, such as images, convey spatial information, while speech data convey dynamic changes in time series. These fundamental differences pose challenges for simultaneous optimization of both modalities, often leading to conflicts during training. For instance, the inclusion of speech data may degrade performance on vision tasks, and vice versa. In addition, traditional speech-to-speech systems rely on separate modules for Automatic Speech Recognition (ASR) and Text-to-Speech, which can increase latency and reduce coherence, limiting their practicality in real-time applications (reddy2023speech, ; fu2024vita, ; zhang2023speechgpt, ; liu2021geometry, ; zhang2025logavatar, ).
@@ 2 Related Work
Recently, thanks to the rapid development of language models such as GPTs (gpt4, ; brown2020language, ), LLaMA (touvron2023llama, ; touvron2023llama2, ), Alpaca (taori2023stanford, ), Vicuna (chiang2023vicuna, ), and Mistral (jiang2023mistral, ), researchers have successfully extended text comprehension to multimodal understanding/reasoning through techniques like multimodal alignment and instruction tuning. For example, models such as LLaVA (liu2023visual, ), Qwen-VL (bai2023qwen, ), Cambrian-1 (tong2024cambrian, ), Mini-Gemini (li2024mini, ), MiniCPM-V 2.5 (hu2024minicpm, ), DeepSeek-VL (lu2024deepseek, ), and SliME (zhang2024beyond, ) have made significant advances in image perception and reasoning, while models like LongVA (zhang2024longva, ) and Video-LLaVA (lin2023video, ) have showcased the latest progress in video understanding. These models are increasingly capable of handling diverse data types, driving the continuous improvement of multimodal perception and understanding capabilities.
Beyond visual modalities, recent years have also witnessed significant progress in incorporating speech capabilities into LLMs, driven by the increasing demand for natural human-computer interaction. The dominant approach has been to cascade ASR, LLM, and TTS modules. This text-centric approach faces fundamental limitations due to the loss of paralinguistic features like tones and emotions. While works like xue2023chat and lin2023paralinguistics have attempted to address these issues by incorporating speech encoders and emotion vectors, they still rely on speech transcription, resulting in substantial latency issues that impact the user experience. The emergence of proprietary models like GPT-4o (NBS_2021, ) has demonstrated the possibility of end-to-end speech interaction, inspiring a new wave of research in speech-enabled MLLMs. Following this trend, several notable works have emerged in the open-source community. Models such as Mini-Omni2 (xie2024mini2, ), LLaMA-Omni (fang2024llama, ), and Moshi (defossez2024moshi, ) have explored various strategies for aligning speech modality with LLMs and achieving duplex dialogue capabilities. While these open-source efforts have successfully enabled duplex speech interaction with LLMs, they still lack the capability to handle visual modalities as demonstrated by GPT-4o, limiting their applications in scenarios requiring both visual and speech understanding.
Despite these advances in both visual and speech modalities, a significant gap remains between proprietary and open-source models. Compared to proprietary models that support multiple modalities, including audio, image, and text, e.g., GPT-4o (gpt4o, ) and Gemini-Pro 1.5 (team2023gemini, ), most open-source models have primarily focused on image and text modalities (zhan2024anygpt, ). Moreover, few open-source models have involved multimodal interaction capabilities, which is a relatively unexplored area. While works like VITA-1.0 (fu2024vita, ) have made initial attempts to introduce speech for human-computer interaction, introducing additional speech data poses challenges to the model’s original multimodal abilities. Furthermore, speech generation typically relies on existing TTS systems, which often results in high latency, thus impacting user experience. In this paper, we present VITA-1.5 that leverages refined training strategies, excelling in perceiving data across four modalities (video, image, text, and audio), while also realizing near real-time vision and speech interaction.
@@
Figure 2: Overall Architecture of VITA-1.5. The input side consists of vision and audio encoders, along with their adapters. The output side has an end-to-end speech generation module, rather than directly using an TTS model.
@@ 3.1 Model Architecture
The overall architecture of VITA-1.5 is depicted in Fig. 2. The input side is the same as that of the VITA-1.0 version (fu2024vita, ), that is, adopting the configuration of “Multimodal Encoder-Adaptor-LLM”. It combines the Vision/Audio Transformer and the Multi-Layer Connector with an LLM for joint training, aiming to enhance the unified understanding of vision, language, and audio. With respect to the output side, VITA-1.5 has its own end-to-end speech module, instead of using the external TTS model like the original VITA-1.0 version.
@@ 3.1.1 Visional Modality
Visual Encoder. VITA-1.5 adopts InternViT-300M [^1] https://huggingface.co/OpenGVLab/InternViT-300M-448px as the visual encoder, with an input image size of 448×448 pixels, generating 256 visual tokens per image. For high-resolution images, VITA-1.5 employs a dynamic patching (chen2024far, ) strategy to capture local details, improving the accuracy of image understanding.
@@ 3.1.2 Audio Modality
Speech Encoder. Similar to (wang2024freeze, ), our audio encoding module consists of multiple downsampling convolutional layers (4x downsampling) and 24 Transformer blocks (with a hidden size of 1024). The downsampling layers help reduce the frame rate of the audio features, improving the processing speed of LLM. The audio encoder has about 350M parameters and an output frame rate of 12.5Hz. Mel-filter bank features are used as the input of the audio encoder, with a window size of 25ms and a shift of 10ms (wang2024freeze, ).
@@
Speech Decoder. TiCodec (ren2024fewer, ) is used as our codec model, customizing a single codebook with a size of 1024. This single-codebook design simplifies the decoding process during the inference phase. The codec model is responsible for encoding continuous speech signals into discrete speech tokens with the frequency of 40Hz, and at the same time has the ability to decode them back into speech signals with the sample rate of 24, [^0] Hz.
The current LLM can only output text tokens, and the speech generation capability requires the LLM to be able to output speech tokens. To this end, we add two speech decoders after the text tokens following (wang2024freeze, ): 1) Non-Autoregressive (NAR) Speech Decoder, which processes text tokens globally and models semantic features, with the aim of generating an initial distribution of speech tokens; 2) Autoregressive (AR) Speech Decoder generates higher quality speech tokens step by step, based on the speech information produced by the NAR decoder. The final sequence of speech tokens is then decoded into a continuous speech signal flow (waveform) using the speech decoder of the Codec model. We adopt 4 LLaMA decoder layers for both NAR and AR speech decoders, where the hidden size is 896 and the parameter size is about 120M.
@@ 3.2 Training Data
Table 1: Training data of multimodal instruction tuning. The images of the synthetic data come from open-source datasets like Wukong (gu2022wukong, ), LAION (schuhmann2022laion, ), and CC12M (changpinyo2021conceptual, ).
@@ 3.3.2 Stage 2: Audio Input Tuning
Stage 2.1 Audio Alignment. After completing the training of Stage 1, the model has developed a strong foundation in image and video understanding. In this stage, our goal is to reduce the discrepancy between audio and language based on Stage 1, enabling the LLM to understand audio inputs. The training data consists of 11,000 hours of speech-transcription pairs. We follow a two-step approach: (a) Speech Encoder Training: We adopt a training framework used in common speech recognition systems, using a Connectionist Temporal Classification (CTC) loss function (graves2006connectionist, ) to train the speech encoder. The aim is for the encoder to predict the transcription text from the speech input. This step ensures that the audio encoder can extract speech features and map them to the text representation space. (b) Speech Adapter Training: After training the speech encoder, we integrate it with the LLM, using an audio adapter to introduce audio features into the input layer of the LLM. The training objective at this stage is to enable the LLM to output the transcription text of the speech data.
@@ 3.3.3 Stage 3: Audio Output Tuning
In the first two stages of training, the VITA-1.5 model has effectively developed its multimodal understanding capabilities. However, a crucial capacity, i.e., speech output, remains absent, which is essential for its role as an interactive assistant. To introduce speech output functionality without compromising the model’s fundamental abilities, we draw on the strategy (wang2024freeze, ), using 3,000 hours of text-speech data and employing a two-step training approach (see Fig. 3).
@@ 4.1 Vision-Language Evaluation
Baselines. We compare a series of open-source MLLMs, including VILA-1.5 (lin2023vila, ), LLaVA-Next (li2024llavanext-strong, ), CogVLM2 (hong2024cogvlm2, ), InternLM-XComposer2.5 (zhang2023internlm, ), Cambrian-1 (tong2024cambrian, ), MiniCPM-V-2.6 (hu2024minicpm, ), Ovis1.5 (lu2024ovis, ), InternVL-Chat-1.5, InternVL-2 (chen2023internvl, ), LLaVA-OV (li2024llava, ), and Video-LLaVA (lin2023video, ), SliME (zhang2024beyond, ), and LongVA (zhang2024longva, ), as well as 5 closed-source MLLMs, including GPT-4V [^7] https://openai.com/index/gpt-4v-system-card/, GPT-4o [^8] https://openai.com/index/hello-gpt-4o/, GPT-4o-mini, Gemini 1.5 Pro (team2023gemini, ), and Claude 3.5 Sonnet [^9] https://www.anthropic.com/news/claude-3-5-sonnet.
Evaluation Benchmarks. To assess the image perception and understanding capabilities of VITA-1.5, we utilize several evaluation benchmarks, including MME (fu2023mme, ), MMBench (liu2023mmbench, ), MMStar (chen2024we, ), MMMU (yue2024mmmu, ), MathVista (lu2023mathvista, ), HallusionBench (guan2024hallusionbench, ), AI2D (hiippala2021ai2d, ), OCRBench (liu2023hidden, ), and MMVet (yu2023mm, ). These benchmarks cover a wide range of aspects, including general multimodal capabilities (e.g., MME, MMBench, and MMMU), mathematical reasoning (MathVista), hallucination detection (HallusionBench), chart (AI2D) and OCR (OCRBench) understanding, providing a comprehensive evaluation results. For video understanding, we use representative evaluation benchmarks including Video-MME (fu2024video, ), MVBench (li2024mvbench, ), and TempCompass (liu2024tempcompass, ).
@@
Mini-Omni2 - - - 4.8 9.8 4.7 9.4
Freeze-Omni 2.8 12.6 14.2 4.2 10.2 4.1 10.5
@@ 4.2 Speech Evaluation
Baselines. The following three baseline models are used for comparison: Wav2vec2-base (baevski2020wav2vec, ), Mini-Omni2 (xie2024mini, ), Freeze-Omni (wang2024freeze, ), and VITA-1.0 (fu2024vita, ).
Evaluation Benchmarks. The Mandarin Evaluation Sets consists of three datasets: aishell-1 (bu2017aishell, ), test net (chen2021gigaspeech, ), and test meeting (zhang2022wenetspeech, ). These datasets are used to evaluate the model’s performance on Mandarin speech. The evaluation metric is the Character Error Rate (CER). The English Evaluation Sets include four datasets: dev-clean, dev-other, test-clean, and test-other (panayotov2015librispeech, ), which are used to evaluate the model’s performance on English speech. The evaluation metric is Word Error Rate (WER). The evaluation results in Table 4 indicate that VITA-1.5 achieves leading accuracy in both Mandarin and English ASR tasks. This demonstrates that VITA-1.5 has successfully integrated advanced speech capability to support multimodal interaction.
5 Conclusion and Future Work
In this paper, we has presented VITA-1.5, a multimodal LLM designed to integrate vision and speech through a carefully crafted three stage training strategy. By relieving the inherent conflicts between modalities, VITA-1.5 achieves robust capabilities in both vision and speech understanding, enabling efficient speech-to-speech interactions without relying on separate ASR or TTS modules. Extensive evaluations demonstrate that VITA-1.5 performs competitively across multimodal benchmarks. We hope that VITA-1.5 can promote the progress of open-source models in the field of real-time multimodal interaction. Although VITA-1.5 has made some contributions, such as multi-modality joint training, end-to-end architecture, response latency, and basic performance, there are two major areas that can be improved in our future work:
1. Personalized MLLM. Currently, VITA-1.5 is generic and do not incorporate individual preferences during interaction. For example, after learning about personal preferences in the interaction, the content and manner of answers can be adjusted accordingly.
2. Long-term memory. The process of human-computer interaction can last 10 minutes or even several hours, in which case it is important for the human-computer interaction in real scenarios.
Impact Statement
This paper studies the technology of large models to enhance their technical level. Its influence is the same as that of the research on other large models and will not be repeated here.
Acknowledgments
This work is funded by National Natural Science Foundation of China (Grant No. 62506158 and No. 62441234), Fundamental Research Funds for the Central Universities, AI & AI for Science Project of Nanjing University (No. 2024300529), and CCF-Tencent Rhino-Bird Open Research Fund.